AB = 3x, BC = 4x, and AC = 42. Find the value of x if B is between A and C. *

Answer by ikleyn(52788)

.

Then 3x + 4x = 42,    7x = 42,   x = 42/7 = 6 units.


Everything is more than OBVIOUS.

Solved, answered and explained.

The complexity level of this assignment is 1 + 2 = 3.